# Improving generalisation via anchor multivariate analysis

### v2

@@ 5 Estimators
$\begin{split}\tilde{\mathbf{X}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{X},\\ \tilde{\mathbf{Y}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{Y},\end{split}$ (21)
where $\Pi_{\mathbf{A}}=\mathbf{A}(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}\in\mathbb{R}^{n\times n}$ projects onto the column space of $\mathbf{A}$, assuming that $\mathbf{A}^{T}\mathbf{A}$ is invertible and $\mathbf{X}$ and $\mathbf{Y}$ are centered.
@@
Finally, build the estimator $\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{\tilde{X}\tilde{Y}})$, where $\mathbf{S}_{\tilde{X}\tilde{Y}}$ is the empirical variance-covariance of the projected data defined in Eq.  (21).
@@
The computational cost of projecting $X$ and $Y$ into the span of $A$ involves computing the covariance matrix of $A$ (of cost $O(nr^{2})$), inverting it (of cost $O(r^{3})$), and two matrix products (of cost $O(nrp)$ and $O(nr^{2})$), resulting in a total complexity of $O(nr^{2}+nrp+r^{3})$. Assuming the unregularised MVA algorithm’s computational cost is $c$, the complexity of its anchor-regularised version is $O(c+nr^{2}+nrp+r^{3})$.
Parameter selection
When knowledge about the graph and anchor variable distributional shift is available, one can estimate the optimal $\gamma$ for robustness against worst-case scenarios. For instance, if the expected anchor variable perturbation strength is up to $1.5$, Rothenhäusler et al. (2018) recommends $\gamma=1.5$. Refer to Section 8.2 of Rothenhäusler et al. (2018) for a detailed estimation approach. Without prior knowledge, various strategies can be considered: Rothenhäusler et al. (2018) propose $\gamma=2$ as a suitable default choice, while Sippel et al. (2021) and Székely et al. (2022) suggest selecting $\gamma$ as a trade-off between prediction error (MSE or $R^{2}$ score) and the correlation between residuals and the anchor variable, or the value of projected residuals in the anchor variable’s span. Similar strategies have been proposed in deemed related algorithms Cortés-Andrés et al. (2022); Li et al. (2022).
Figure 2: Robustness to increasing perturbation strength for PA, unregularised, IV, anchor-regularised ($\gamma=5$) algorithms ( MLR, OPLS, RRR, PLS and CCA). Anchor versions show robustness in terms of $R^{2}$ for bounded intervention strength.
High Dimensional Estimators
Dealing with high dimensional data challenges estimators and often requires introducing regularisation Bühlmann and Van De Geer (2011); Candes and Tao (2007). Regularisation has been generally introduced to control models’ capacity and avoid overfitting Girosi et al. (1995); Tibshirani (1996); Hastie et al. (2009), but also to introduce prior (domain) knowledge in the algorithms, as in AR. Adding a regularisation term $\Omega(\|\mathbf{\Theta}\|)$ to the empirical loss in Eq. (20) leads to estimate $\mathbf{\Theta}$ as
$\begin{split}\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}&(\mathbf{S}_{XY})\\ +(\gamma-1)&f_{\mathbf{\Theta}}(\mathbf{S}_{XY|A})+\Omega(\|\mathbf{\Theta}\|).\end{split}$ (22)
In CCA, OPLS, and PLS regression (or the rank in RRR), the number of components can be viewed as regularisation hyperparameters. Consequently, the optimisation task for anchor-regularised MVA algorithms may involve many hyperparameters. For instance, in anchor-regularised Reduced Rank Ridge Regression (A-RRRR), three hyperparameters require tuning: $\ell_{2}$ regularisation $\alpha$, rank $\rho$, and anchor regularisation $\gamma$. Since each hyperparameter optimization addresses different objectives, it may be impractical to optimise all of them simultaneously. Sippel et al. (2021) advise selecting hyperparameters as a trade-off between prediction performance (e.g., MSE or $R^{2}$) and a proxy of robustness to anchor intervention (e.g. correlation between anchor and residuals).
6 Experiments
Our theoretical findings are substantiated through an extensive series of experiments that highlight the robustness of anchor-regularised MVA algorithms to increasing perturbation strength (cf. Section 6.1). Furthermore, we provide a concrete high-dimensional example to elucidate the process of hyperparameter selection (cf. Section 6.1), and demonstrate its practical relevance by showcasing how anchor regularisation enhances climate predictions in a real-world Climate Science application (cf. Section 6.2).
6.1 Simulation Experiments
Robustness to Perturbation Strength
To demonstrate how anchor regularisation can be applied with different multivariate analysis algorithms, we adapt the experiments from (Rothenhäusler et al., 2018, Sec. 2.2) to a multioutput setting.
We assume that the training data $(A,X,Y)$ follows a distribution based on the linear SCM:
$\begin{split}\varepsilon_{A},\varepsilon_{H},&\varepsilon_{X},\varepsilon_{Y}\sim\mathcal{N}(0,1)\\ A&\leftarrow\varepsilon_{A}\\ H&\leftarrow\varepsilon_{H}\\ X&\leftarrow A+H+\varepsilon_{X}\\ Y&\leftarrow\mathbf{W}^{T}X+H+\varepsilon_{Y},\end{split}$ (23)
where $X$ and $Y$ are multivariate, of respective dimension $d=p=10$ ($d=p=300$ with $n=200$ in the high dimensional setting), and $\mathbf{C}$ is low rank (see Supplementary  10.1 for more details on the experiments settings). We generate test data by intervening on the anchor variable’s distribution, setting $A\sim\mathcal{N}(0,t)$, where $t$ is the perturbation strength. Each MVA algorithm assumes oracle knowledge of the rank $\rho$ of $\mathbf{C}$, aligning RRR’s rank, PLS regression’s component count, and CCA accordingly. IV setting ($A$ only affects $Y$ through $X$) and we show that even in this scenario, anchor regularised algorithms (with $\gamma=5$) outperform the PA, IV, and unregularised algorithms.
Shown in Fig. 2, all anchor-compatible algorithms exhibit robustness to distribution shifts. Anchor-regularised models (with $\gamma=5$) excel with bounded-strength interventions; PA regularisation is optimal for weak interventions; and IV regularisation for unlimited perturbation strength. Overall, anchor-regularised algorithms maintain stable performance across various perturbation strengths.
We also conducted a set of high-dimensional experiments ($p\gg n$ and $d\gg n$) using RRR and MLR both regularised with $\ell_{2}$ norm (ridge regularisation). In both cases, the anchor-regularised algorithms exhibit robustness to increasing perturbation strength (see Fig. 6 in §10.1). This is also the case when $A$ is a confounder (affecting both $X$ and $Y$, potentially through $H$), in which case anchor-regularised algorithms present are equally optimal for a wide range of perturbation strength (see Fig. 8 in §10.1). Though not as pronounced as in anchor-compatible algorithms, anchor-regularised CCA also displays robustness to perturbations in $A$, prompting further investigation into its behaviour.
Hyperparameter Selection
We showcase how hyperparameter selection can be performed through a simulation experiment in a high-dimensional setting using an anchor-regularised version of Reduced Rank Ridge Regression (Mukherjee and Zhu, 2011).
In this setting, we are given three hyperparameters: $\gamma$ enforces robustness to intervention on the anchor, $\alpha$ and $\rho$ aim to maximise prediction performance. Following (Sippel et al., 2021) we select hyperparameters as a trade-off between predictive performance (measured via MSE) and correlation between anchor and residuals. We give equal weights to both objectives but this can be adapted regarding the knowledge available and the application. Thus hyperparameters are selected at testing such that they minimise the combination of the two objectives. Fig. 3 shows how augmenting regularisation for distributional robustness to anchor intervention often reduces predictive performance in the training sample (Fig. 3.A), yet the trade-off translates into improved predictive performance in testing samples. More details about this experiment setting are in  §10.2.

### v3

@@ Appendix B Proofs
$\begin{split}\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f(C_{XY})]&=f(\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T})+\gamma f(\mathbf{D}\mathbf{M}\Sigma_{A}\mathbf{M}^{T}\mathbf{D}^{T}).\end{split}$ (18)
Let’s now note that for $(A,X,Y)\sim\mathbb{P}_{\text{train}}$ we have
$\operatorname{\mathbb{E}_{\text{train}}}\left[\begin{pmatrix}X\\ Y\end{pmatrix}|A\right]$ $=\mathbf{D}\operatorname{\mathbb{E}_{\text{train}}}[\varepsilon|A]+\mathbf{D}\mathbf{M}\operatorname{\mathbb{E}_{\text{train}}}[A|A]$
$=\mathbf{D}\mathbf{M}A,$
where $\mathbb{P}_{\text{train}}$ is the training distribution and $\operatorname{\mathbb{E}_{\text{train}}}[.]=E_{(X,Y)\sim\mathbb{P}_{\text{train}}}[.]$ the expectation with regard to $\mathbb{P}_{\text{train}}$.
As $\varepsilon$ is mean centred and independent of $A$. Thus, we can write
$f(C_{XY|A})=f(\mathbf{D}\mathbf{M}C_{A}\mathbf{M}^{T}\mathbf{D}^{T})$
by linearity of $f$. Taking its expectation over the training distribution, we get
$\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY|A})]=f(\mathbf{D}\mathbf{M}\Sigma_{A}\mathbf{M}^{T}\mathbf{D}^{T}),$ (19)
which is similar to the right term of Eq. (18) as $\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY|A})]=f(\Sigma_{XY|A})$. A similar reasoning leads to
$\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY})]-\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY|A})]$ $=f(\mathbf{D}\Sigma_{\varepsilon}\mathbf{D}^{T}).$ (20)
Pluging-in Eq. (19) and Eq. (20) in Eq. (18) we get
$\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f(C_{XY})]$ $=\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY})]+(\gamma-1)\operatorname{\mathbb{E}_{\text{train}}}[f(C_{XY|A})]$
$=f(\Sigma_{XY})+(\gamma-1)f(\Sigma_{XY|A}),$
which concludes the proof. ∎
B.2 Proof of Proposition
Proof.
The loss can easily decomposed as
$\mathbb{E}[f_{\mathbf{\Theta}}(R(\operatorname{\mathbf{B}^{\diamond}})\otimes R(\operatorname{\mathbf{B}^{\diamond}}))]$ $=\mathbb{E}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))]+f_{\mathbf{\Theta}}((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T})$
$-\mathbb{E}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\begin{pmatrix}X\\ Y\\ H\end{pmatrix})]$
$\geq\mathbb{E}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))]+f_{\mathbf{\Theta}}((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T})$
$-\mathbb{E}[f_{\mathbf{\Theta}}((R(\mathbf{B})\otimes R(\mathbf{B})))]^{\frac{1}{2}}\mathbb{E}[f_{\mathbf{\Theta}}(((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T}))^{\frac{1}{2}}$
Where the first equality is a simple bias-variance decomposition and we use Cauchy-Schartz inequality for the inequality. By assumptions
$\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\operatorname{\mathbf{B}^{\diamond}})\otimes R(\operatorname{\mathbf{B}^{\diamond}}))]\leq\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))].$
This leads to
$\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))]$ $\geq\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))]+f_{\mathbf{\Theta}}((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}^{do(A:=\nu)}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T})$
$-\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}((R(\mathbf{B})\otimes R(\mathbf{B})))]^{\frac{1}{2}}f_{\mathbf{\Theta}}(((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}^{do(A:=\nu)}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T}))^{\frac{1}{2}}$
which is equivalent to
$f_{\mathbf{\Theta}}((\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}^{do(A:=\nu)}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T})$ $\leq 4\mathbb{E}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))]$
$\leq 4f_{\mathbf{\Theta}}(\Sigma_{\varepsilon})+4\gamma f_{\mathbf{\Theta}}(\mathbf{M}\Sigma_{A}\mathbf{M}^{T}).$
This concludes the proof. ∎
B.3 Proof of Equivalence for Transformed Data
Proof.
From
$\begin{split}\tilde{\mathbf{X}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{X}\\ \tilde{\mathbf{Y}}&=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{Y},\end{split}$ (21)
we can easily derive that
$\mathbf{S}_{\tilde{X}\tilde{Y}}=$ $\frac{1}{n-1}(\tilde{\mathbf{X}}\tilde{\mathbf{Y}})^{T}(\tilde{\mathbf{X}}\tilde{\mathbf{Y}})$
$=$ $\frac{1}{n-1}((\mathbf{X}\mathbf{Y}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}}\mathbf{X}\mathbf{Y})^{T}(\mathbf{X}\mathbf{Y}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}}\mathbf{X}\mathbf{Y}))$
As $\Pi_{\mathbf{A}}$ is an orthogonal projection, we have that $(\mathbf{X}\mathbf{Y})^{T}(\Pi_{\mathbf{A}}\mathbf{X}\mathbf{Y})=\Pi_{\mathbf{A}}(\mathbf{X}\mathbf{Y})^{T}\mathbf{X}\mathbf{Y}=\mathbf{S}_{XY|A}$, leading to
$\mathbf{S}_{\tilde{X}\tilde{Y}}=$ $\mathbf{S}_{XY}+2(\sqrt{\gamma}-1)\mathbf{S}_{XY|A}+(\sqrt{\gamma}-1)^{2}\mathbf{S}_{XY|A}$
$=$ $\mathbf{S}_{XY}+(\gamma-1)\mathbf{S}_{XY|A}.$
Thus, by linearity of $f_{\mathbf{\Theta}}$, the minimiser of $f_{\mathbf{\Theta}}(\mathbf{S}_{\tilde{X}\tilde{Y}})$ is given by
$\mathbf{\Theta}^{\gamma}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{\tilde{X}\tilde{Y}})$
$=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{XY})+(\gamma-1)f_{\mathbf{\Theta}}(\mathbf{S}_{XY|A}).$
This conclude the proof. ∎
Appendix C ESTIMATORS: FURTHER DETAILS
Simpler Formulation
For computational and practical reasons, $\widehat{\mathbf{\Theta}}^{\gamma}$ can be estimated by transforming the training data:
$\tilde{\mathbf{X}}$ $=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{X}$ (22)
$\tilde{\mathbf{Y}}$ $=(\mathbf{I}+(\sqrt{\gamma}-1)\Pi_{\mathbf{A}})\mathbf{Y},$ (23)
@@ Appendix C ESTIMATORS: FURTHER DETAILS
where $\Pi_{\mathbf{A}}=\mathbf{A}(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}\in\mathbb{R}^{n\times n}$ projects onto the column space of $\mathbf{A}$, assuming that $\mathbf{A}^{T}\mathbf{A}$ is invertible and $\mathbf{X}$ and $\mathbf{Y}$ are centered. Finally, build the estimator $\widehat{\mathbf{\Theta}}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\mathbf{S}_{\tilde{X}\tilde{Y}})$, where $\mathbf{S}_{\tilde{X}\tilde{Y}}$ is the empirical variance-covariance of the projected data defined in Eq.  (22). Proof of the equivalence of using transformed data as in Eq. (22) is equivalent to Eq. (11) is available in Supplementary  B.3. We give computational complexity and consistency results in Supplementary C.
@@
The computational cost of projecting $X$ and $Y$ into the span of $A$ involves computing the covariance matrix of $A$ (of cost $O(nr^{2})$), inverting it (of cost $O(r^{3})$), and two matrix products (of cost $O(nrp)$ and $O(nr^{2})$), resulting in a total complexity of $O(nr^{2}+nrp+r^{3})$. Assuming the unregularised MVA algorithm’s computational cost is $c$, the complexity of its anchor-regularised version is $O(c+nr^{2}+nrp+r^{3})$. We note that this generally incurs an affordable additive computational cost, as it is equivalent to two linear regressions. The only main concern arises with high-dimensional $A$, which is not addressed in this work.
Appendix D SIMULATION EXPERIMENTS: FURTHER DETAILS
D.1 Experiment Setting
The results of the simulation experiments are obtained as follows. The training data are sampled from the following SCM
$\begin{split}\varepsilon_{A},\varepsilon_{H},&\varepsilon_{X},\varepsilon_{Y}\sim\mathcal{N}(0,1)\\ A&\leftarrow\varepsilon_{A}\\ H&\leftarrow\varepsilon_{H}\\ X&\leftarrow A\mathbf{1}_{p}^{\top}+H\mathbf{1}_{p}^{\top}+\varepsilon_{X}\\ Y&\leftarrow\mathbf{W}^{T}X+H\mathbf{1}_{d}^{\top}+\varepsilon_{Y},\end{split}$ (24)
and the testing data are generated by modifying $\varepsilon_{A}$ such that $\varepsilon_{A}\sim\mathcal{N}(0,t)$. The perturbation strength $t$ is varied over a linear sequence $t\in[0,4]$ with $20$ steps. We repeat each experiment $B=10$ times by sampling $n=300$ training and testing samples of $(A,X,Y)$. We plot the average Mean Squared Errors in Fig. 2 and Fig. 7. Both $A$ and $H$ are one-dimensional, while $X$ and $Y$ are $10$-dimensional. The matrix $\mathbf{W}$ is generated as a low-rank (of rank $\rho$) matrix such that $\mathbf{W}=\mathbf{U}\mathbf{V}$ with $\mathbf{A}\in\mathbb{R}^{d\times\rho}$ and $\mathbf{B}\in\mathbb{R}^{\rho\times p}$. The matrices $\mathbf{\Gamma}_{i}$ are hThe coefficients of $\mathbf{A}$ and $\mathbf{B}$ are sampled uniformly between $1$ and $2$ and are normalised such that their sum is $1$.
We employ the algorithms CCA, PLS, and MLR implemented in the scikit-learn library (BSD 3-Clause License) Pedregosa et al. (2011), respectively CCA, PLSRegression and LinearRegression learners. We use the code available at https://github.com/rockNroll87q/RRRR (MIT License)  Svanera et al. (2019); Mukherjee and Zhu (2011) for the Reduce Rank Regression algorithm and our implementation of OPLS based on Arenas-García and Gómez-Verdejo (2015) using an eigenvalue decomposition as we use the constraint $A^{T}A=\mathbf{I}$.
High-Dimensional Setting
We also conduct a high-dimensional experiment to evaluate the performance of anchor-regularized algorithms when the dimensionality of $X$ and $Y$ exceeds the sample size. We generate data by sampling $n=200$ instances of $X$ and $Y$, each with a dimensionality of $d=p=300$. The rank of $\mathbf{W}$ is set to $\rho=100$. We compare the results between Ridge Regression and Reduced Rank Ridge Regression. As shown in Figure 5, both methods demonstrate robustness across a wide range of perturbation strengths.
Figure 5: Experiments in high-dimensional setting with $d=p=300$ and $n=200$. We can see that both Multi-output Ridge Regression and Reduced rank Ridge Regression are optimal for a wide range of perturbation strength. Shaded areas represent the range of two standard errors of the mean for running $B=20$ times the experiment.
Non-Gaussian Noise Experiments
To assess the robustness of our results presented in paragraph 6.1, we conducted the same toy model experiments with noise $\varepsilon_{A}$, $\varepsilon_{H}$, $\varepsilon_{X}$, and $\varepsilon_{Y}$ following an exponential distribution (Fig. 7.A) and a gamma distribution (Fig. 7.B) with scales $1$.
As observed in Fig. 7, anchor-regularised models remain optimal for a wide range of perturbation strengths, except for anchor-regularised CCA. The behavior of anchor-regularised CCA is explained by its incompatibility with anchor regularisation.
Confounding Anchor Experiment
We also reproduced experiments from equation 12 in Rothenhäusler et al. (2018) to demonstrate how anchor-regularised MVA algorithms exhibit interesting robustness properties when the anchor has a confounding effect. These results are illustrated in Fig. 8. Here, the training distribution is encapsulated in the following Structural Causal Model (SCM):
$\begin{split}\varepsilon_{A},\,\varepsilon_{H},\,&\varepsilon_{X},\,\varepsilon_{Y}\sim\mathcal{N}(0,1),\\ A&\leftarrow\varepsilon_{A},\\ H&\leftarrow A+\varepsilon_{H},\\ X&\leftarrow A+H+\varepsilon_{X},\\ Y&\leftarrow\mathbf{W}^{T}X+H+\varepsilon_{Y},\end{split}$ (25)
with again testing distribution sampled by setting $\varepsilon_{A}\sim\mathcal{N}(0,t)$ where $t$ is the perturbation strength and $\mathbf{W}$ being lower rank.